# Peierls substitution and Hall motion in exotic Carroll dynamics

### v1

@@ Abstract
The particle with first-order dynamics proposed by Dunne, Jackiw and Trugenberger (DJT) to justify the “Peierls substitution” is obtained by reduction from both of two-parameter centrally extended Galilean and Carroll systems. In the latter case the extension parameters $\kappa_{exo}$ and $\kappa_{mag}$ generate non-commutativity of the coordinates resp. behave as an internal magnetic field. The position and momentum follow uncoupled anomalous Hall motions. Consistently with partial immobility, one of the Carroll boost generators is broken but the other remains a symmetry. Switching off $\kappa_{exo}$, the immobility of unextended Carroll particles is recovered. The Carroll system is dual to an uncharged anyon on the horizon of a black hole which exhibits the spin-Hall effect.
@@ I Introduction
The fundamental property which has long delayed interest in Carrollian physics is that particles with Carroll symmetry can not move Leblond ; SenGupta . In this paper we show that a particle associated with the recently discovered “exotic” two-parameter central extension of the planar Carroll group Azcarraga ; Ancille0 ; Ancille1 ; Ancille2 ; Marsot21 analogous to the one of the planar Galilei group in LLGal does move, namely by following the Hall law. However we confirm that turning off the exotic extension a massive Carroll particle does indeed not move.
@@
Carrollian physics has become lately a much discussed topic Gibbons02 ; Carrollvs ; Bergshoeff14 ; Morand1 ; Morand2 ; Marsot21 ; deBoer ; deBoer23 , justified by its applications to gravitational waves, dark matter and to black hole physics, see Carroll4GW ; Laura19 ; Laura22 ; Laura23 ; MZH ; Gray ; Aggarwal ; Ecker23 ; Ecker24 ; MZCH and many further references. Our new results in this paper complete previous work Leblond ; deBoer ; Casalbuoni23 ; MultiCarroll .
We start with recalling the Peierls substitution proposed many years ago in many-body physics Peierls ; Luttinger ; Kohn ; Azbel ; Hofstadter  : for a charged particle in the plane subjected to perpendicular magnetic and electric fields and constrained to stay in the lowest Landau level (i) the components of quantum position operator should be non-commuting and (ii) its Hamiltonian should reduce to the mere potential and have no kinetic term,
@@
In DHPeierls we re-derived the DJT system from the non-commutative mechanics associated with the exotic (double) extension of the planar Galilei group noticed by Lévy-Leblond LLGal and further investigated in Kosinski ; Grigore ; LSZ ; DHPeierls ; DHJPA ; NCLandau ; ExoRev ; ZHChiral ; PlyushChiral1 ; PlyushChiral2 by Hamiltonian reduction FaTMF ; FaJa . Our new paper carries out a similar study in the hitherto unexplored exo-Carrollian case, recalling occasionally the Galilean one for the sake of comparison.
We proceed in three steps. After recalling the DJT system (sec.II) we show that it can in fact be derived from both exotic Galilei and exotic Carroll particles. Then focusing our attention at the latter, we show that a charged exotic Carroll particle moves by following an anomalous Hall law (IV.7) Ezawa ; Stone both in the presence, but also in the absence of an external magnetic field, reminiscent of the anomalous Hall effect Kar+ ; AnomHall . Turning off the exotic extension we recover an usual Carroll particle with its celebrated immobility.
For simplicity, we limit our attention to constant electromagnetic fields $B,\,{\mathbf{E}}$ in the plane. Since we work in flat euclidean space, vectors will have lower indices.
@@ II The Dunne-Jackiw-Trugenberger (DJT) system
Legendre transformation of (II.1) then yields the Poisson brackets and Hamiltonian,
$\Big{\{}{\xi}_{i},{\xi}_{j}\Big{\}}=-\frac{1}{eB\,}\epsilon_{ij}{\quad\text{and}\quad}{\mathscr{H}}=eV({\bm{\xi}})\,,$ (II.3)
@@
  • The Hall law (II.2) implies partial immobility: the particle moves perpendicularly to the electric field, ${\mathbf{E}}=-{\bm{\nabla}}V$. When the electric field vanishes, ${\mathbf{E}}=0$, then the DJT particle does not move at all, ${\bm{\xi}}={\bm{\xi}}_{0}=\mathop{\rm const.}\nolimits$
    to be compared with the usual circular motions of a charged massive particle in a magnetic field.
We note for later use that that the motions can also be obtained by finding the kernel of the Souriau 2-form SSD ; Marsot21 ,
@@
see the Appendix A.
In the next section we show tht the DJT system can be obtained by reduction from two different, centrally extended particle models. After recalling some relevant facts from the Galilean case for comparison, we carry out a similar but more detailed analysis for Carroll.
@@ III Exotic Galilean dynamics
As shown in DHPeierls ; DHJPA ; NCLandau ; ExoRev ; ZHChiral , the DJT system is obtained from “exotic” Galilean mechanics Kosinski ; Grigore ; LSZ ; PlyushChiral1 ; PlyushChiral2 associated with the two-parameter central extension of the planar Galilei group LLGal by Hamiltonian reduction FaTMF ; FaJa . One of the parameters present in any dimension is the mass, $m$ Bargmann . The other, “exotic” one we denote here by $\kappa_{exo}$ is specific for planar physics LLGal . In terms of the non-commutativity parameter and the effective mass,
$\theta=\frac{\kappa_{exo}}{m^{2}}{\quad\text{and}\quad}m^{*}=m(1-e{\theta}{B})\,,$ (III.1)
@@
$\{x_{i},x_{j}\}=\frac{\theta}{1-e{\theta}{B}}\,\epsilon_{ij},\quad\{x_{i},p_{j}\}=\frac{1}{1-e{\theta}{B}}\,\delta_{ij},\quad\{p_{i},p_{j}\}=\frac{eB}{1-e{\theta}{B}}\,\epsilon_{ij}\,$ (III.4)
or alternatively, from the kernel of the exotic Souriau form $\sigma=\Omega-d{\mathscr{H}}\wedge dt$ cf. (A.4),
@@
$p_{1}(t)=\cos\left(\frac{eB}{m^{*}}t\right)\,c_{3}-\sin\left(\frac{eB}{m^{*}}t\right)\,c_{4}+\frac{m\,E_{2}}{B}\,,$ (III.8)
$p_{2}(t)=-\cos\left(\frac{eB}{m^{*}}t\right)\,c_{4}-\sin\left(\frac{eB}{m^{*}}t\right)\,c_{3}-\frac{m\,E_{1}}{B}\,.$ (III.9)
The arbitrary constants $c_{i}$ fix the initial position, and the radii of the circles.
It was shown in earlier work DHPeierls ; DHJPA ; NCLandau ; ExoRev that when the magnetic field takes the critical value
@@
or equivalently when the effective mass vanishes, $m^{*}=0$, we have  :
@@
In the critical case (III.10) the momentum is determined by the field. Both the position coordinate ${\bm{x}}$ moves according to the Hall law.
@@
More insight is obtained by combining the position and the momentum into the guiding centre Ezawa ; Stone ; ZHChiral ,
@@
When $m^{*}\neq 0$, then the guiding centre moves, independently of the non-commutativity parameter $\theta$ and even outside the critical value (III.10), according to the Hall law,
@@
The trajectories shown in FIG. 1 are circles which drift along the guiding centre which itself follows the Hall law (III.13) consistently with FIG.s #3 and #4 of ZHChiral .
Figure 1: (i) Exotic Galilean motions with parameters $e=1,\,B=1,\,\bm{E}=(0,1),\,m=1,\,c_{1}=c_{2}=c_{3}=0,\,c_{4}=3$. The blue/purple curves are below ($\theta=0.5$) / above ($\theta=2$) the critical value $\theta=1$. The orange curve is for large $|\theta|$ ($\theta=-2$). (ii) When $\theta\to 1$ either from the left or from the right, the frequency diverges and changes sign.
When the critical case (III.10) is approached, the rotation speeds up and its angular velocity $\omega^{*}={eB}/{m^{*}}$ diverges. Crossing the critical value the direction of the rotation is reversed.
For $|\theta|$ very large instead the effective mass diverges, $m^{*}\to\infty$, the frequency tends to 00 and the motion reduces to the (displaced) guiding centre,
@@
Further progress is achieved by rewriting the Souriau form (III.5) in terms of $Q_{i}$,
$\sigma_{exoG}=\sigma_{QG}+\sigma_{pG}=\qquad$ (III.15)
[…]
where the first describes the (Hall) dynamics of the guiding centre (III.13), and $\sigma_{pG}$ determines that of the momentum which however intervenes in real-space dynamics.
When the magnetic field takes the critical value in (III.10) or equivalently, when the effective mass vanishes, $m^{*}=0$, then Hamiltonian reduction shows DHJPA ; NCLandau that the phase 4-dimensional space reduces to a 2-dimensional one upon which the guiding centre (III.12) is a good coordinate. The reduced Hamiltonian structure
$\Omega_{red}=\frac{1}{2\theta}d{\bf Q}\times d{\bf Q}\quad\;\Leftrightarrow\quad\;\Big{\{}Q_{1},Q_{2}\Big{\}}=-\theta=-\frac{1}{eB_{crit}}\,,\quad{\mathscr{H}}_{red}=eV(\mathbf{Q})\,$ (III.16)
@@
${\mathscr{L}}_{red}=\frac{1}{2\theta}\mathbf{Q}\times{\mathbf{Q}}^{\prime}-V(\mathbf{Q})\,,$ (III.17)
Comparing with the DJT expressions (II.4) and (II.1), we recover the DJT system in section II with the guiding centre ${\bf Q}$ replacing ${\bm{\xi}}$.
For later comparison, the reduced Souriau form is
$\sigma_{exoG}^{red}=\hbox{$\textstyle\frac{1}{2}$}eB\epsilon_{ij}dQ_{i}\wedge dQ_{j}+\,eE_{i}dQ_{i}\wedge dt.$ (III.18)
@@ IV Exotic Carroll particle
A remarkable fact which has long escaped attention and was recognised only recently Azcarraga ; Ancille0 ; Ancille1 ; Ancille2 ; Marsot21 is that, unlike in $d\geq 3$ dimensions, the planar Carroll group admits also a two-parameter central extension The associated particle will be called an exotic Carroll particle. (Our notations correspond to $\kappa_{exo}=m^{2}\theta=2q_{2},\;\kappa_{mag}=-2q_{1}$ in Marsot21 .) We assume that our particle has non-vanishing mass and electric charge, $m\neq 0$ and $e\neq 0$, respectively, and is subjected to constant planar magnetic and electric fields. Unlike as for Galilei Bargmann , the mass $m$ is for Carroll an externally given constant and not a central-extension parameter.
In terms of the non-commutativity parameter and the effective magnetic field,
@@
respectively, Souriau’s orbit construction SSD applied to the exotic Carroll group yields the Souriau form Ancille0 ; Ancille1 ; Ancille2 ; Marsot21 ,
$\sigma_{exoC}=\underbrace{d{\bm{p}}\wedge d{\bm{x}}+\frac{{B^{*}}}{2}\,\epsilon_{ij}dx_{i}\wedge dx_{j}+\frac{\theta}{2}\,\epsilon_{ij}dp_{i}\wedge dp_{j}}_{\Omega_{exo}}+eE_{i}\,dx_{i}\wedge ds\,.$ (IV.2)
Comparison with (III.5) shows that (up to $t\leadsto s$ and $eB\leadsto B^{*}$) this is identical to the Galilean expression except for the missing kinetic term in the Hamiltonian which is merely the “naked” potential,
${{\mathscr{H}}}_{Car}=eV\,.$ (IV.3)
@@
${\mathscr{L}}_{Car}={\bm{p}}\cdot\dot{{\bm{x}}}+\hbox{$\textstyle\frac{1}{2}$}B^{*}\epsilon_{ij}x_{i}\dot{x}_{j}+\hbox{$\textstyle\frac{1}{2}$}\theta\,\epsilon_{ij}p_{i}\dot{p}_{j}-eV\,,$ (IV.4)
where the “dot” means derivation w.r.t. Carroll time, $d/ds$. The absence of the kinetic term implies that the Carroll equations of motion are decoupled for all values of the parameters Marsot21 ,
@@
Then introducing again the effective mass,
@@
When $m^{*}\neq 0$ the system is regular and the coordinates follow an anomalous Hall law Ezawa ; Stone ,
$\dot{x}_{i}=-\big{(}\frac{e\theta}{1-\theta B^{*}}\big{)}\,\epsilon_{ij}E_{j}\;{\quad\Rightarrow\quad}\;x_{i}(s)=-\big{(}\frac{e\theta}{1-\theta B^{*}}\big{)}\,\epsilon_{ij}{E_{j}}s+x_{i}^{0}\,,$ (IV.7)
supplemented by
$\dot{p}_{i}=\big{(}\frac{1}{1-\theta B^{*}}\big{)}eE_{i}\;{\quad\Rightarrow\quad}\;\;p_{i}(s)=\big{(}\frac{eE_{i}}{1-\theta B^{*}}\big{)}\,s+p^{0}(s)\,.\qquad\;$ (IV.8)
@@
The $p$-equation has no effect on the dynamics of the position, though, and could actually be ignored when studying the ${\bm{x}}$-dynamics. We recall for comparison that Galilean particles follow the Hall law (III.11) only in the critical case (III.10) when the effective mass vanishes.
@@
Figure 2: (i) Hall motions for various values of the non-commutativity parameter $\theta\neq 1$, after fixing the other parameters as $e=1,\,B^{*}=1,\,{\mathbf{E}}=(0,1)$ so that the motion remains perpendicular to the electric field ${\mathbf{E}}$. For $\theta=0$ we recover the unextended Carroll situation with no motion. (ii) When $\theta\to 1$ the velocity diverges. For very large $\theta$ instead, the residual velocity tends to the Hall value $\dot{x}_{1}=eE/B^{*}$ (to be compared with the Galilean case in FIG.1).
@@
$m^{*}=0\quad\Leftrightarrow\quad\left({\theta}{B^{*}}\right)_{crit}={1}\,,$ (IV.9)
the velocity diverges, cf. FIG.2(ii). The coefficient in (IV.7) changes sign when the denominator crosses zero, flipping over the direction of the motion, as in FIG.2ii. — to be compared with the reversal of the rotation in exotic Galilean mechanics ZHChiral in FIG.1. The novelty is that we get Hall-like Carroll motion always, not only in the critical case.
An interesting particular case is obtained by presenting (IV.7) as,
$\big{(}\underbrace{1-\kappa_{mag}\theta}_{0}-e\theta{B}\big{)}\dot{x}_{i}=-e\theta\epsilon_{ij}E_{j}\,.$ (IV.10)
@@
$\theta\kappa_{mag}=1\quad\Leftrightarrow\quad\kappa_{exo}\kappa_{mag}=m^{2}\,.$ (IV.11)
Then the exotic factor $e\theta$ drops out (as long as it does not vanish), yielding :
@@
Eqn. (IV.10) suggests to consider yet another remarkable particular case:

### v3

@@ Abstract
The particle with first-order dynamics proposed by Dunne, Jackiw and Trugenberger (DJT) to justify the “Peierls substitution” is obtained by reduction from both of the planar two-parameter centrally extended Galilean and Carroll systems. In the latter case the extension parameters $\kappa_{exo}$ and $\kappa_{mag}$ generate non-commutativity of the coordinates resp. behave as an internal magnetic field. The position and momentum follow uncoupled anomalous Hall motions. Consistently with partial immobility, one of the Carroll boost generators is broken but the other remains a symmetry. Switching off $\kappa_{exo}$, the immobility of unextended Carroll particles is recovered. The Carroll system is dual to an uncharged anyon on the horizon of a black hole which exhibits the spin-Hall effect. Physical applications are shortly reviewed.
@@ I Introduction
The fundamental property which has long delayed interest in Carrollian physics is that particles with Carroll symmetry can not move Leblond ; SenGupta . In this paper we show that a particle associated with the recently discovered “exotic” two-parameter central extension of the planar Carroll group Azcarraga ; Ancille0 ; Ancille1 ; Ancille2 ; Marsot21 analogous to the one of the planar Galilei group LLGal does move, namely by following the Hall law. However we confirm that turning off the exotic extension a massive Carroll particle does indeed not move.
@@
Carrollian physics has become lately a much discussed topic Gibbons02 ; Carrollvs ; Bergshoeff14 ; Morand1 ; Morand2 ; Marsot21 ; deBoer ; deBoer23 , justified by its applications to gravitational waves, dark matter and to black hole physics, see Carroll4GW ; Laura19 ; Laura22 ; Laura23 ; MZH ; Gray ; Aggarwal ; Ecker23 ; Ecker24 ; MZCH and many further references.
This paper completes previous work on Carroll dynamics Leblond ; deBoer ; Casalbuoni23 ; Marsot21 ; MZH ; MultiCarroll . We start with recalling the Peierls substitution proposed many years ago and widely used in many-body physics Peierls ; Luttinger ; Kohn ; Azbel ; Hofstadter  : for a charged particle in the plane subjected to perpendicular magnetic and electric fields and constrained to stay in the lowest Landau level (i) the components of quantum position operator should be non-commuting and (ii) its Hamiltonian no kinetic term and reduce to the mere potential,
@@
Some time ago DHPeierls ; DHJPA we re-derived the DJT system from the non-commutative mechanics associated with the “exotic” (double) extension of the planar Galilei group noticed by Lévy-Leblond LLGal and further investigated by Hamiltonian reduction Kosinski ; Grigore ; LSZ ; DHPeierls ; DHJPA ; NCLandau ; ExoRev ; ZHChiral ; PlyushChiral1 ; PlyushChiral2 ; FaTMF ; DHJPA . In this paper we carry out a similar study in the hitherto unexplored exotic Carrollian case, recalling the Galilean one for the sake of comparison. For simplicity, we limit our attention to constant electromagnetic fields, $B\neq 0$
We proceed as follows. After recalling the DJT system (sec.II) we show that it can in fact be derived from both exotic Galilei and exotic Carroll particle models by Faddeev-Jackiw reduction. Focusing our attention henceforth at the second case, we show that a charged exotic Carroll particle moves by following an anomalous Hall law (IV.7) Ezawa ; Stone both in the presence, but also in the absence of an external magnetic field, reminiscent of the anomalous Hall effect Kar+ ; AnomHall . Turning off the exotic extension we recover the usual Carroll particle with its celebrated immobility.
The behaviour of the various physical systems is related to their curious symmetry properties. In the DJT case the Hall field breaks the Galilei boost symmetry. However for the same system, the Carroll boost generator along the electric field is a symmetry while the perpendicular one is broken: we have what we call a “half-Carroll” symmetry.
The same behaviour is found for exotic particles put into a Hall field: Galilei boosts are fully broken for Galilei particles, however half-Carroll symmetry is found for Carroll particles  [^5] Full Carroll symmetry could be restored in an “enlarged” framework.. The symmetry behaviour of DJT is confirmed by reduction of the exotic-particle cases. An outlook at previous and ongoing research is given in Sec.VII.
@@ II The Dunne-Jackiw-Trugenberger (DJT) system
Legendre transformation of (II.1) then yields the Poisson brackets
$\Big{\{}{\xi}_{i},{\xi}_{j}\Big{\}}=-\frac{1}{eB\,}\epsilon_{ij}{\quad\text{and}\quad}{\mathscr{H}}_{DJT}=eV({\bm{\xi}})\,,$ (II.3)
@@
  • The Hall law (II.2) implies partial immobility: the particle moves perpendicularly to the electric field, ${\mathbf{E}}=-{\bm{\nabla}}V$. When the electric field vanishes, ${\mathbf{E}}=0$, then the DJT particle does not move, ${\bm{\xi}}={\bm{\xi}}_{0}=\mathop{\rm const.}\nolimits$ – in contrast with the circular motions of a usual charged massive particle in a magnetic field.
We note for later use that the motions can also be obtained by finding the kernel of the Souriau 2-form SSD ; Marsot21
@@
see Appendix A.
In the next section we derive the DJT system by reduction from two different, centrally extended particle models. After recalling some relevant facts from the Galilean case for comparison, we carry out a similar but more detailed analysis for Carroll.
@@ III Exotic Galilean dynamics
The planar Galilei group admits a two-parameter central extension LLGal leading to “exotic” Galilean mechanics Kosinski ; Grigore ; LSZ ; PlyushChiral1 ; PlyushChiral2 One of the parameters, present in any dimension, is the mass, $m$ Bargmann . The other, “exotic” one we denote here by $\kappa_{exo}$ is specific for planar physics LLGal and implies the non-commutativity of the coordinates DHPeierls ; DHJPA ; NCLandau ; ExoRev ; ZHChiral . In terms of the non-commutativity parameter and the effective mass,
$\theta=\frac{\kappa_{exo}}{m^{2}}{\quad\text{and}\quad}m^{*}=m(1-{\theta}{eB})\,,$ (III.1)
@@
$\{x_{i},x_{j}\}=\frac{\theta}{1-{\theta}{eB}}\,\epsilon_{ij},\quad\{x_{i},p_{j}\}=\frac{1}{1-{\theta}{eB}}\,\delta_{ij},\quad\{p_{i},p_{j}\}=\frac{eB}{1-{\theta}{eB}}\,\epsilon_{ij}\,$ (III.4)
or alternatively, from the kernel of the exotic Souriau form $\sigma=\Omega-d{\mathscr{H}}\wedge dt$ (A.5),
@@
$p_{1}(t)=\;\;\;c_{3}\,\cos\left(\frac{eB}{m^{*}}t\right)-c_{4}\,\sin\left(\frac{eB}{m^{*}}t\right)+\frac{m\,E_{2}}{B}\,,$ (III.8)
$p_{2}(t)=\;\,-c_{4}\,\cos\left(\frac{eB}{m^{*}}t\right)-c_{3}\,\sin\left(\frac{eB}{m^{*}}t\right)-\frac{m\,E_{1}}{B}\,.$ (III.9)
The arbitrary constants $c_{i}$ fix the initial position and the radii of the circles.
When the effective mass vanishes, $m^{*}=0$, which happens when the magnetic field takes the critical value
@@
then we have DHPeierls ; DHJPA ; NCLandau ; ExoRev  :
@@
In the critical case (III.10) the momentum is determined by the field. The position coordinate ${\bm{x}}$ moves according to the Hall law.
@@
Further insight is obtained by combining the position and the momentum into the guiding centre Johnson:1950zz ; Ezawa ; Stone ; ZHChiral ,
@@
When $m^{*}\neq 0$, then the guiding centre moves, independently of the non-commutativity parameter $\theta$ (and not only the critical value (III.10)), according to the Hall law,
@@
as illustrated by FIG.1 .
Figure 1: Exotic Galilean motions with parameters $\,m=1,\,e=1,\,B=1,\,\bm{E}=(0,1),\,c_{1}=c_{2}=c_{3}=0,\,c_{4}=3$. The blue/purple curves are below ($\theta=0.5$) / above ($\theta=2$) the critical value $\theta=1$. The orange curve is for large $|\theta|$ ($\theta=-2$).
For $|\theta|$ very large instead the effective mass diverges, $m^{*}\to\infty$, the frequency tends to 00 and the real-space motion reduces to the (displaced) guiding centre (III.12),
@@
When the critical case (III.10) is approached, the rotation speeds up and its angular velocity $\omega^{*}={eB}/{m^{*}}$ diverges, FIG.2. Crossing the critical value the direction of the rotation is reversed.
Figure 2: When $eB\theta\to 1$ either from the left or from the right, the frequency diverges and changes sign.
When the magnetic field takes the critical value in (III.10) or equivalently, when the effective mass vanishes, $m^{*}=0$, then Hamiltonian reduction FaTMF ; FaJa shows DHJPA ; NCLandau that the phase 4-dimensional space reduces to a 2-dimensional one for which the guiding centre $Q_{i}$ in (III.12) is a good coordinate. The reduced Hamiltonian structure
$\Omega^{red}=\frac{1}{2\theta}d{\bf Q}\times d{\bf Q}\quad\;\Leftrightarrow\quad\;\Big{\{}Q_{1},Q_{2}\Big{\}}=-\theta=-\frac{1}{eB_{crit}}\,,\quad{\mathscr{H}}^{red}=eV(\mathbf{Q})\,$ (III.15)
@@
${\mathscr{L}}^{red}=\frac{1}{2\theta}\mathbf{Q}\times{\mathbf{Q}}^{\prime}-V(\mathbf{Q})\,.$ (III.16)
The reduced Souriau form is
$\sigma_{G}^{red}=\hbox{$\textstyle\frac{1}{2}$}eB_{crit}\epsilon_{ij}dQ_{i}\wedge dQ_{j}+\,eE_{i}dQ_{i}\wedge dt\,.$ (III.17)
Comparing these results with (II.1), (II.3), (II.4) and (II.5) we recognize the DJT system with the guiding centre ${\bf Q}$ replacing ${\bm{\xi}}$. In conclusion,
Proposition III.4.
The DJT system is the Hamiltonian reduction of the exotic Galilei particles.
@@ IV Exotic Carroll particle
A remarkable fact recognised only recently Azcarraga ; Ancille0 ; Ancille1 ; Ancille2 ; Marsot21 is that, unlike in $d\geq 3$ dimensions, the planar Carroll group admits also a two-parameter central extension. The associated particle will be called an exotic Carroll particle. (Our notations correspond to $\kappa_{exo}=m^{2}\theta=2q_{2},\;\kappa_{mag}=-2q_{1}$ in Marsot21 ).
We assume that our particle has non-vanishing mass and electric charge, $m\neq 0$ and $e\neq 0$, respectively. Unlike as for Galilei Bargmann , the mass $m$ is for Carroll an externally given constant and not a central-extension parameter. In terms of the non-commutativity parameter and the effective magnetic field,
@@
respectively, Souriau’s orbit construction SSD applied to the exotic Carroll group yields the Souriau form Ancille0 ; Ancille1 ; Ancille2 ; Marsot21 . Assuming that the particle is subjected to constant Hall-type electromagnetic field, the Souriau form
$\sigma_{exoC}=\underbrace{d{\bm{p}}\wedge d{\bm{x}}+\frac{\theta}{2}\,\epsilon_{ij}dp_{i}\wedge dp_{j}+\frac{{B^{*}}}{2}\,\epsilon_{ij}dx_{i}\wedge dx_{j}}_{\Omega_{exo}}\;+\;eE_{i}\,dx_{i}\wedge ds\,.$ (IV.2)
Comparison with (III.5) shows that (up to $t\leadsto s$ and $eB\leadsto B^{*}$) this is identical to the Galilean expression except for the missing kinetic term in the Hamiltonian which, for Carroll, is merely the “naked” potential,
${{\mathscr{H}}}_{Carr}=eV\,.$ (IV.3)
@@
${\mathscr{L}}_{Carr}={\bm{p}}\cdot\dot{{\bm{x}}}+\hbox{$\textstyle\frac{1}{2}$}B^{*}\epsilon_{ij}x_{i}\dot{x}_{j}+\hbox{$\textstyle\frac{1}{2}$}\theta\,\epsilon_{ij}p_{i}\dot{p}_{j}-eV\,,$ (IV.4)
where the “dot” means derivation w.r.t. Carroll time, $d/ds$. The absence of the kinetic term implies that the Carroll equations of motion are fully decoupled for all values of the parameters Marsot21 ,
@@
Then introducing again the effective mass analog of (III.1),
@@
When $m^{*}\neq 0$ the system is regular and the coordinates follow, for all $\theta$ and $B^{*}$, an anomalous Hall law Ezawa ; Stone
$\dot{x}_{i}=-\big{(}\frac{e\theta}{1-\theta B^{*}}\big{)}\,\epsilon_{ij}E_{j}\,,$ (IV.7)
@@
Figure 3: (i) Hall motions for various values of the non-commutativity parameter $\theta\neq 1$, after fixing the other parameters as $e=1,\,B^{*}=1,\,{\mathbf{E}}=(0,1)$ so that the motion remains perpendicular to the electric field ${\mathbf{E}}$. For $\theta=0$ we recover the unextended Carroll situation with no motion. (ii) When $\theta\to 1$ the velocity diverges. For very large $\theta$ instead, the residual velocity tends to the Hall value $\dot{x}_{1}=eE/B^{*}$ (to be compared with FIG.2 in the Galilei case).
Eqn. (IV.7) is actually supplemented by
$\dot{p}_{i}=\big{(}\frac{1}{1-\theta B^{*}}\big{)}eE_{i}\,,$ (IV.8)
which however has no effect on the position.
@@
$m^{*}=0\quad\Leftrightarrow\quad{\theta}{B^{*}}_{crit}={1}\,,$ (IV.9)
the velocity diverges, cf. FIG.3(ii). The coefficient in (IV.7) changes sign when the denominator crosses zero, flipping over the direction of the motion, as in FIG.3ii. — to be compared with the reversal of the rotation in exotic Galilean mechanics ZHChiral in FIG.1. For comparison, Galilean particles follow the Hall law only in the critical case (III.10) when the effective mass vanishes.
$\bullet$ An interesting particular case is obtained by presenting (IV.7) as,
$\big{(}\underbrace{1-\kappa_{mag}\theta}_{0}-{\theta}{eB}\big{)}\dot{x}_{i}=-e\theta\epsilon_{ij}E_{j}\,.$ (IV.10)
@@
$\theta\kappa_{mag}=1\quad\Leftrightarrow\quad\kappa_{exo}\kappa_{mag}=m^{2}\,,$ (IV.11)
the exotic factor $e\theta$ drops out (as long as it does not vanish), yielding :
@@
For $\theta=0$ we would recover the no-motion property $\dot{x}_{i}=0$.
For simplicity, we assume henceforth that $m=1$.
$\bullet$ Eqn. (IV.10) suggests to consider yet another remarkable particular case: